OpenStudy (princessaurora):

**Will fan & Medal** A survey of factories in five northeastern states found that 10% of the 300 workers surveyed were satisfied with the benefits offered by their employers. The margin of error for the population proportion is ± BLANK %. What is the blank?

OpenStudy (princessaurora):

please help me set up the problem

OpenStudy (anonymous):

what's the sample size?

OpenStudy (princessaurora):

I don't know

OpenStudy (princessaurora):

i need help finding it

OpenStudy (anonymous):

try reading the first sentence and see what you come up with

OpenStudy (princessaurora):

10/300=.03?

OpenStudy (anonymous):

what does sample size mean to you? remember, a survey has been conducted and there were 300 respondents. 10 of the 300 respondents were satisfied with the benefits offered by their employers.

OpenStudy (princessaurora):

the 300 people are the sample size?

OpenStudy (anonymous):

yes, very good. in the formula\[MoE = z^*\sqrt{\frac{\hat{p} \left( 1-\hat{p} \right)}{n}}\]n is the sample size. MoE is the Margin of Error.

OpenStudy (anonymous):

what is the sample proportion? (this is \(\hat{p}\) in the formula)

OpenStudy (princessaurora):

is it 10% or 10/300=0.3?

OpenStudy (anonymous):

\(\hat{p}\) is proportion... it can be given directly as a proportion or percentage, or calculated as the number (out of the total satisfying a given requirement) divided by the total. hint: 10/300 is not appropriate and is it the percentage divided by the total. .10*300 = 30 (which is 10% of 300) is the number of respondents who were satisfied by the benefits offered by their employers.

OpenStudy (princessaurora):

so p is 30?

OpenStudy (anonymous):

proportion means ratio... percentage is also a ratio but with 100 as the denominator (always!) to calculate the proportion of respondents who were satisfied with the benefits offered by their employers, we need the number of respondents who were satisfied with the benefits offered by their employers (call this x) AND the total number of respondents (the sample size). \(\hat{p}=\Large{\frac{x}{n}}\) in your problem, they've given it to you... it's the only percentage mentioned.

OpenStudy (princessaurora):

so 10%?

OpenStudy (anonymous):

yep. next decide what value for \(z^*\) you'd leik to use. please check the second link I provided to see what value you think is appropriate

OpenStudy (anonymous):

btw: you must use the decimal form of the percentage in the formula

OpenStudy (princessaurora):

so .10 for 10%?

OpenStudy (princessaurora):

Now I'm trying to find z I need help

OpenStudy (anonymous):

yes, .10 for 10% from the link: Note: when you see a margin of error in a news report, it it almost always referring to a 95% confidence interval. But other levels of confidence are possible. since your problem doesn't give a level of confidence, follow the suggestion above and use the appropriate \(z^*\) value.

OpenStudy (anonymous):

the link has a table, relating confidence levels and \(z^*\) values

OpenStudy (anonymous):

you could also use a standard normal table or your calculator

OpenStudy (princessaurora):

so does z have no value? do i just use z?

OpenStudy (anonymous):

no, did you see table 10.1 in the link?

OpenStudy (princessaurora):

yeah i don't understand what z stands for in the table

OpenStudy (anonymous):

well, forget that for just a second and let's make sure you get the correct value. what level of confidence should you use (according to the site and common practice)?

OpenStudy (anonymous):

from the link: Note: when you see a margin of error in a news report, it it almost always referring to a 95% confidence interval. But other levels of confidence are possible. since your problem doesn't give a level of confidence, follow the suggestion above.

OpenStudy (princessaurora):

2.0 or more precisely 1.96?

OpenStudy (princessaurora):

i wasnt paying attention and completely missed what you said the confidence level was

OpenStudy (princessaurora):

wow im dumb sorry

OpenStudy (princessaurora):

you're an angel for putting up with my slow self

OpenStudy (anonymous):

yes, you should use a 95% level of confidence, corresponding to a \(z^*\) value of 1.96 this 1.96 is how many standard deviations from the mean, in a normal distribution.

OpenStudy (anonymous):